Graph both of equations (y=3x+1 and 2y-6x=6)on a graph. Is the pair parallel or perpendicular?
Answer by mananth(16946)

y=3x+1..........................(1)
2y-6x=6
/2
y-3x=3
y=3x+3..........................(2)
Find the x & y intercept points by plugging x=0,y=0 in both equations.
equation I
when x= 0 y= -3 ,( 0 , 1 )
when y = 0 x= 3 ,( -1/3 , 0 )
Equation 2
when x= 0 y= 3 ( 0 , 3 )
when y = 0 x= 1 ( -1 , 0 )





The lines are parallel